OpenStudy (anonymous):

The reciprocal of 4plus the reciprocal of 5 is the reciprocal of what number?

OpenStudy (anonymous):

(1/4) + (1/5) = 9/20, which is the reciprocal of 20/9

OpenStudy (anonymous):

Thanks! [:

OpenStudy (anonymous):

1/4+1/5=5/20+4/20 =9/20